# Similarity-Dissimilarity Loss for Multi-label Supervised Contrastive Learning

### v4

@@ Appendix E More Experiment Results: Evaluation on Medical Domain
Appendix F Additional Discussion of Related Work
@@
Multi-label classification presents significant challenges due to its inherent label correlations, extreme and sparse label spaces, and long-tailed distributions. For instance, in the International Classification of Diseases (ICD) [34, 35], the presence of one label (e.g., "Pneumococcal pneumonia") may increase the probability of co-occurring labels (e.g., "fever" or "cough"). Furthermore, multi-label datasets frequently exhibit long-tailed distributions, where a small subset of labels occurs with high frequency while the majority appear rarely. This imbalance typically results in models that perform adequately on common labels but underperform on infrequent ones [7, 36, 37]. Additionally, the number of potential label combinations increases exponentially with the number of labels, resulting in heightened computational complexity and substantial memory requirements.
@@
Contrastive learning aims to learn a representation of data such that similar instances are close together in the representation space, while dissimilar instances are far apart. Compared to self-supervised contrastive learning, such as SimCLR [5] and MoCo [6], Khosla et al. [1] proposed supervised contrastive learning, which fully leverages class annotation information to enhance representations within the contrastive learning framework. Recent studies have extended supervised contrastive learning from single-label to multi-label scenarios by exploiting the additional information inherent in multi-label tasks. Zhang et al. [2] proposed a hierarchical multi-label representation learning framework specifically designed to utilize comprehensive label information while preserving hierarchical inter-class relationships.
@@
In subsequent research, Zhang and Wu [4] developed Multi-Label Supervised Contrastive Learning (MulSupCon), featuring a novel contrastive objective function that expands the positive sample set based on label overlap proportions. Similarly, the Jaccard Similarity Probability Contrastive Loss (JSPCL) [3] employed the Jaccard coefficient [38] to calculate label similarity between instances, sharing conceptual foundations with MulSupCon [4] and MSC loss [10] that those approaches primarily focus on similarity only, but ignoring dissimilarity.
@@
Despite these advancements, the intricate relationships and dependencies between multi-label samples have yet to be fully elucidated. To address this gap, we introduce multi-label relations and formalize the concepts of similarity and dissimilarity. Inspired by the idea of re-weighting of logit adjustment [39], focal loss [9] and class-balanced loss [40], we leverage the similarity and dissimilarity factors to re-weight the contrastive loss, thereby enhancing discriminative power in multi-label scenarios.

### v5

@@ G.3 Evaluation on Medical Domain
Table 4 reports that our method achieves superior results on MIMIC-IV-ICD9-Full compared to MIMIC-III-Full, despite both datasets employing identical ICD-9 coding standards. This marked performance differential can be attributed primarily to the more extensive training corpus available in MIMIC-IV-ICD9-Full (see in Table 2). While MIMIC-IV-ICD10-Full similarly comprises a substantial volume of clinical data, its considerably expanded label taxonomy introduces increased representational sparsity and presents additional computational and methodological challenges (Nguyen et al., 2023). Moreover, the MIMIC-IV-ICD10-50 dataset demonstrates consistent performance metrics in this restricted setting, providing empirical evidence that label space dimensionality constitutes a critical determinant of model training efficacy.
Comparative analysis of model performance reveals that Llama significantly outperforms RoBERTa across evaluation metrics, a finding attributable to scaling laws of LLMs and the extensive knowledge and training corpus during the pre-training phase (Kaplan et al., 2020; Bahri et al., 2024; Huang et al., 2024b). Although LLMs demonstrate considerable efficacy in domain-specific applications (Nori et al., 2023), our results indicate that PLM-ICD consistently surpasses both RoBERTa and Llama across all experimental configurations. This hierarchical performance pattern aligns with theoretical expectations, as PLM-ICD incorporates architecture and training paradigms specifically optimized for automated ICD coding tasks (Huang et al., 2022). Despite the increasing generalization capabilities of foundation models in diverse applications, significant questions persist regarding their capacity to achieve state-of-the-art performance on highly specialized tasks, particularly within the medical domain, without substantial domain-specific training or parameter-efficient adaptation techniques (Saab et al., 2024). Contemporary research on foundation model applications in biomedical domain has predominantly relied on specialized adaptation methods tailored to specific domain requirements. The comparative advantages of domain-specific pre-training becomes particularly evident following the development of initial foundation model architectures, as exemplified by widely implemented medical models such as Med-PaLM (Singhal et al., 2023) and Med-Gemini(Saab et al., 2024).
Table 6: Comparison with previous state-of-the-art methods. MRR and AKIL rely on DRG codes, CPT codes, and medications, which are manually annotated for each sample by human coders. These methods are included for reference, although a direct comparison with our approach is not strictly fair. Our method adopts the MSCL framework with Similarity–Dissimilarity Loss and a PLM-ICD encoder.
Models AUC F1 P@8
Macro Micro Macro Micro
CAML (Mullenbach et al., 2018) 89.5 98.6 8.8 53.8 70.9
MSATT-KG (Xie et al., 2019) 91.0 99.2 9.0 55.3 72.8
MSMN (Yuan et al., 2022) 95.0 99.2 9.4 58.4 75.2
KEPTLongformer (Yang et al., 2022) - - 11.8 59.9 77.1
PLM-ICD (Huang et al., 2022) 92.6 98.9 10.4 59.8 77.1
PLM-CA (Edin et al., 2024) 91.6 98.9 8.9 57.3 74.5
CoRelation (Luo et al., 2024) 95.2 99.2 10.2 59.2 76.2
GKI-ICD (Zhang et al., 2025) 96.2 99.3 12.3 62.3 77.7
Our 94.5 99.4 12.5 62.3 78.4
MRR (Wang et al., 2024b) 94.9 99.5 11.4 60.3 77.5
AKIL (Wang et al., 2024c) 94.8 99.4 11.2 60.6 78.4
Appendix H Additional Discussion of Related Work
@@
Multi-label classification presents significant challenges due to its inherent label correlations, extreme and sparse label spaces, and long-tailed distributions. For instance, in the International Classification of Diseases (ICD) (Edin et al., 2023; Ji et al., 2024), the presence of one label (e.g., ”Pneumococcal pneumonia”) may increase the probability of co-occurring labels (e.g., ”fever” or ”cough”). Furthermore, multi-label datasets frequently exhibit long-tailed distributions, where a small subset of labels occurs with high frequency while the majority appear rarely. This imbalance typically results in models that perform adequately on common labels but underperform on infrequent ones (Zhang et al., 2023; Huang et al., 2023; Wang et al., 2024a). Additionally, the number of potential label combinations increases exponentially with the number of labels, resulting in heightened computational complexity and substantial memory requirements.
@@
Contrastive learning aims to learn a representation of data such that similar instances are close together in the representation space, while dissimilar instances are far apart. Compared to self-supervised contrastive learning, such as SimCLR (Chen et al., 2020) and MoCo (He et al., 2020), Khosla et al. (2020) proposed supervised contrastive learning, which fully leverages class annotation information to enhance representations within the contrastive learning framework. Recent studies have extended supervised contrastive learning from single-label to multi-label scenarios by exploiting the additional information inherent in multi-label tasks. Zhang et al. (2022) proposed a hierarchical multi-label representation learning framework specifically designed to utilize comprehensive label information while preserving hierarchical inter-class relationships.
@@
In subsequent research, Zhang & Wu (2024) developed Multi-Label Supervised Contrastive Learning (MulSupCon), featuring a novel contrastive objective function that expands the positive sample set based on label overlap proportions. Similarly, the Jaccard Similarity Probability Contrastive Loss (JSPCL) (Lin et al., 2023) employed the Jaccard coefficient (Jaccard, 1912) to calculate label similarity between instances, sharing conceptual foundations with MulSupCon (Zhang & Wu, 2024) and MSC loss (Audibert et al., 2024) that those approaches primarily focus on similarity only, but ignoring dissimilarity.
@@
Despite these advancements, the intricate relationships and dependencies between multi-label samples have yet to be fully elucidated. To address this gap, we introduce multi-label relations and formalize the concepts of similarity and dissimilarity. Inspired by the idea of re-weighting of logit adjustment (Menon et al., 2021), focal loss (Lin et al., 2017) and class-balanced loss (Cui et al., 2019), we leverage the similarity and dissimilarity factors to re-weight the contrastive loss, thereby enhancing discriminative power in multi-label scenarios.
Appendix I Limitations
Our experimental analysis on the PASCAL dataset reveals a crucial insight: the significant advantages of the proposed method appear primarily in multi-label dense scenarios. On PASCAL, where the average number of labels per instance is only approximately $1.5$, the observed performance improvements are marginal compared to datasets such as MS-COCO or NUS-WIDE, which exhibit richer multi-label structures. This outcome is consistent with the theoretical underpinnings of our approach.
The limited gains can be attributed to the structural characteristics of PASCAL. With relatively few labels per instance, the task effectively reduces to a near single-label classification problem, particularly under constrained batch sizes. As demonstrated in Equation 17, under such conditions the proposed Similarity–Dissimilarity Loss degenerates to the standard supervised contrastive loss, thereby diminishing the benefit of re-weighting.
More broadly, in datasets with low label cardinality, the scarcity of relational diversity limits the expressiveness of our framework. Consequently, the advantages of dynamically re-weighted contrastive objectives become less pronounced, underscoring the dependence of our method on datasets that exhibit complex and diverse multi-label interactions.
Appendix J The Use of LLMs Statement
We use LLMs for spelling, grammar and typo checking.